The mass of proton is 1.0073 u and that of neutron is 1.0087 u (u = atomic mass unit). The binding energy of 2He4is


        (a)    0.061 u


(b)    0.0305 j


        (c)    0.0305 erg


        (d)    28.4 MeV


Grade:10

1 Answers

Aditi Chauhan
10 years ago

(d)

Δm = (2 × 1.0074 + 2 × 1.0087 - 4.0015) = 0.0307

E = (Δm) × 931 Mev = 0.0307 × 931 = 28.5 Mev